OpenStudy (anonymous):

Factor: 12m^3-8m^2=15m

hartnn (hartnn):

subtract 15m from both sides first, what u get ?

hartnn (hartnn):

:O

hartnn (hartnn):

the question asks to factor, so you'll stop at m(6m + 5)(2m - 3) = 0

OpenStudy (nurali):

12m^3 - 8m^2 - 15m = 0 m(12m^2 - 8m - 15) = 0 m(12m^2 - 18m + 10m - 15) = 0 m((12m^2 - 18m) + (10m - 15)) = 0 m(6m(2m - 3) + 5(2m - 3)) = 0 m(6m + 5)(2m - 3) = 0